# Non-convolutional Graph Neural Networks

### v1

@@ 1 Introduction: Convolutions in GNNs
Graph neural networks (GNNs) [1, 2, 3, 4, 5]—neural models operating on representations of nodes ($\mathcal{V}$) and edges ($\mathcal{E}$) in a graph (denoted by $\mathcal{G}=\{\mathcal{V},\mathcal{E}\},\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$, with structure represented by the adjacency matrix $A_{ij}=\mathbbm{1}[(v_{i},v_{j})\in\mathcal{E}]$)—have shown promises in a wide array of social and physical [6, 7, 8, 9, 10] modeling applications. Most GNNs follow a convolutional scheme, where the $D$-dimensional node representations $\mathbf{X}\in\mathbb{R}^{|V|\times D}$ are aggregated based on the structure of local neighborhoods:
@@
In this paper, we propose a variant of GNN that does not engage the convolution operator (Equation 1) at all. [^1] Code at: https://anonymous.4open.science/r/rum-834D/ Specifically, we stochastically sample a random walk terminating at each node and use the anonymous experiment associated with the random walk to describe its topological environment. This is combined with the semantic representations along the walk and fed into a recurrent neural network layer [21] to form the node embedding, which we call the unifying memory. We theoretically (Section 4) and experimentally (Section 5) show that the resulting model, termed random walk with unifying memory (RUM) relieves the aforementioned symptoms and offers a compelling alternative to the popular convolution-based GNNs.
@@ 2 Related works: ways to walk on a graph
More importantly, focusing mostly on expressiveness, no aforementioned walk-based or random walk kernel-based GNNs address the issue of over-smoothing and over-squashing in GNNs. Some of these works are also de facto convolutional, as the random walks are only incorporated as features in the message-passing scheme. Interestingly, most of these works are either not applicable to, or have not been tested on, node-level tasks. In the experimental Section 5, we show that RUM not only outperforms these baselines on most graph-level tasks (Table 3) but also competes with a wide array of state-of-the-art convolutional GNNs on node-level tests (Table 2). Moreover, random walk kernel-based architectures, which explicitly enumerate random walk-generated subgraphs are typically much slower than WL-equivalent convolutional GNNs, whereas RUM is faster than even the simplest variation of convolutional GNN (Figure 4).
@@ 3 Architecture: combining topologic and semantic trajectories of walks
where $D(v_{i})=\sum A_{ij}$ is the degree of the node $v_{i}$. Walks originating from or terminating at any given node $v$ can thus be easily generated using this Markov chain. We record the trajectory of embeddings associated with the walk as $\omega_{x}(w)=(\mathbf{X}_{i})=(\mathbf{X}_{0},\mathbf{X}_{1},\ldots,\mathbf{X}_{l}).$ In this paper, we only consider finite-long $l$-step random walk $|w|=l\in\mathbb{Z}^{+}$. In our implementation, the random walks are sampled ad hoc during each training and inference step directly on GPU using Deep Graph Library [41] (see Appendix Section A). Moreover, the walk considered here is not necessarily a path, as repeated traversal of the same node $v_{i}=v_{j},i\neq j$ is not only permitted, but also crucial to effective topological representation, as discussed below.
@@
Obvious choices to model $f$ include a feed-forward neural network after concatenation, and $\phi_{x},\phi_{u}$ recurrent neural networks (RNNs). This implies that, different from most convolutional GNNs, parameter sharing is natural and the number of parameters is going to stay constant as the model incorporates a larger neighborhood. Compared to dot product-based, transformer-like modules [43], RNNs not only have linear complexity (see detailed discussions below) w.r.t. the sequence length but also naturally encodes the inductive bias that nodes closer to the origin have stronger impact on the representation. The gated recurrent unit (GRU)[21] variant is used everywhere in this paper. Additional regularizations are described in Appendix Section B.1.
@@
Assumption 2 is easy to satisfy for feed-forward assuneural networks [45] and RNNs [46]. Composing injective functions, we remark that $h(w)$ is also injective w.r.t. $\omega_{x}(w)$ and $\omega_{u}(w)$; despite of Assumption 3, our analysis can be extended to disjointed graphs by restricting the analysis to the connected regions in a graph. Under such assumptions, we show, in Remark 8 (deferred to the Appendix), that $\psi$ is injective, meaning that nodes with different random walks will have different distributions of representations $\psi(v_{1})\neq\psi(v_{2})$. We also refer the readers to the Theorem 1 in Micali and Zhu [42] for a discussion on the representation power of anonymous experiments on unlabelled graphs. Combining with the semantic representations and promoting the argument from a node level to a graph level, we arrive at:
@@
The main idea of the proof of Theorem 4 (in Appendix Section D.2) involves explicitly enumerating all possible non-isomorphic structures for graphs with 3 nodes and showing, by induction, that if the theorem stands for graph of $N-1$ size it also holds for $N$-sized graphs. We also show in Appendix Section D.1 that a number of key graph properties such as cycle size (Example 8.1) and radius (Example 8.2) that convolutional GNNs struggle [48, 49] to learn can be analytically expressed using $\omega_{u}$. As these are solely functions of $\omega_{x}$, they can be approximated arbitrarily well by universal approximators. These examples are special cases of the finding that RUM is stricly more expressive than Weisfeiler-Lehman isomorphism test [14]:
@@ 4.2 RUM alleviates over-smoothing and over-squashing
is deferred to Appenxix Section B.2, where we verify that the inter-node Jacobian $|\partial\mathbf{X}_{v}^{(l+1)}/\partial\mathbf{X}_{u}^{(0)}|$ decays slower as the distance between $u,v$ grows vis-à-vis the convolutional counterparts. Briefly, although RUM does not address the information bottleneck with exponentially growing reception field (the $1/(\hat{A}^{l+1})_{uv}$ term in Equation 16), it nevertheless can have a non-vanishing (nor exploding) gradient from the aggregation function ($|\nabla\phi_{x}|$).
@@ 5 Experiments
On a wide array of real-world node- and graph-level tasks, we benchmark the performance of RUM to show its utility in social and physical modeling. Next, to thoroughly examine the performance of RUM, we challenge it with carefully designed illustrative experiments. Specifically, we ask the following questions in this section, with Q1, Q2, and Q3 already theoretically answered in Section 4: Q1: Is RUM more expressive than convolutional GNNs? Q2: Does RUM alleviate over-smoothing? Q3: Does RUM alleviate over-squashing? Q4: Is RUM slower with convolutional GNNs? Q5: Is RUM robust? Q6: How does RUM scale up to huge graphs? Q7: What components of RUM are contributing most to the performance of RUM?
@@
Having illustrated in Section 4 that RUM can distinguish non-isomorphic graphs, we experimentally test this insight on the popular Circular Skip Link dataset [63, 64]. Containing 4-regular graph with edges connected to form a cycle and containing skip-links between nodes, this dataset is artificially synthesized in Murphy et al. [63] to create an especially challenging task for GNNs. As shown in Appendix Table 7, all convolutional GNNs fail to perform better than a constant baseline (there are 10 classes uniformly distributed). 3WLGNN [65], a higher-order GNN of at least $\mathcal{O}(2)$ complexity that operates on explicitly enumerated triplets of graphs, can distinguish these highly similar 4-regular graphs by comparing subgraphs. RUM, with $\mathcal{O}(|\mathcal{N}|)$ linear complexity, achieves similarly high accuracy. One can think of RUM as a stochastic approximation of the higher-order GNN, with all of its explicitly enumerated subgraphs being identified by RUM with a probability that decreases with the complexity of the subgraph.
@@
Having studied the relationship between inference speed and the walk lengths and number of samples, we furthermore study its impact on performance. Using Cora [50] citation graph and vary the walk lengths and number of samples from $1$ to $9$, where the performance of RUM improves as more samples are taken and longer walks are employed, though more than 4 samples and walks longer than $L>4$ yield qualitatively satisfactory results; this empirical finding has guided our hyperparameter design. In Figure 2, we also compare the Dirichlet energy of RUM-generated layer representations with not only baselines GCN [1] and GAT [11], but also strategies to alleviate over-smoothing discussed in Section 2, namely residual connection and stochastic regularization [40, 31, 66], and show that when $L$ gets large, only RUM can maintain Dirichlet energy. Traditionally, since Kipf and Welling [1] (see its Figure 5 compared to Figure 4), the best performance on Cora graph was found with $2$ or $3$ message-passing rounds, since mostly local interactions are dominating the classification, and more rounds of message-passing almost always lead to worse performance. As theoretically demonstrated in Section 4.2, RUM is not as affected by these symptoms. Thus, RUM is especially appropriate for modeling long-range interactions in graphs without sacrificing local representation power.
@@
To verify that RUM indeed alleviates over-squashing (Section B.2), in Figure 6, we adopt the tree neighborhood matching synthetic task in Alon and Yahav [19] where binary tree graphs are proposed with the label and the attributes of the root matching a faraway leave. The full discussion is moved to the Appendix Section B.3.
@@
Though both have linear runtime complexity (See Section 3), intuitively, it might seem that RUM would be slower than convolution-based GCN due to the requirement of multiple random walk samples. This is indeed true for CPU implementations shown in Figure 4 left. When equipped with GPUs (specs in Appendix Section A), however, RUM is significantly faster than even the simplest convolutional GNN—GCN [1]. It is worth mentioning that the GCN and GAT [11] results were harvested using the heavily optimized Deep Graph Library [41] sparse implementation whereas RUM is implemented naïvely in PyTorch [67], though the popular GRU component [21] have already undergone CUDA-specific optimization.
@@ 6 Conclusions
We design an innovative GNN that uses an RNN to unify the semantic and topological representations along stochastically sampled random walks, termed random walk with unifying memory (RUM) neural networks. Free of the convolutional operators, our methodology does not suffer from symptoms characteristic of Laplacian smoothing, including limited expressiveness, over-smoothing, and over-squashing. Most notably, our method is more expressive than the Weisfeiler-Lehman isomorphism test and can distinguish all non-isomorphic graphs up to the reconstruction conjecture. Thus, it is more expressive than all of convolutional GNNs equivalent to the WL test, as we demonstrate theoretically in Section 4. RUM is significantly faster on GPUs than even the simplest convolutional GNNs (Section 5) and shows superior performance across a plethora of node- and graph-level benchmarks.
@@
Very dense graphs. As evidenced by the underwhelming performance of the Computer [54] dataset and discussed in Section 5, RUM might suffer from high variance with very dense graphs (average degree over 15). Tottering. In our implementation, we have not ruled out the 2-cycles from the random walks, as that would require specialized implementation for walk generation. This, however, would reduce the average information content in a fixed-length random walk (known as tottering [69]). We plan to characterize the effect of excluding these walks. Biased walk. Here, we have only considered unbiased random walk, whereas biased random walk might display more intriguing properties as they tend to explore faraway neighborhoods more effectively [22]. Directed graphs. Since we have only developed RUM for undirected graph (random walk up to a random length is not guaranteed to exist for directed graphs), we would have to symmetrize the graph to perform on directed graphs (such as the heterophilic datasets [55]); this create additional information loss and complexity.
@@
Theoretical. We plan to expand our theoretical framework to account for the change in layer width and depth to derive analytical estimates for realizing key graph properties. Applications. Random walks are intrinsically applicable to uncertainty-aware learning. We plan to incorporate the uncertainty naturally afforded by the model to design active learning models. On the other hand, the physics-based graph modeling field is also heavily dominated by convolutional GNNs. Inspired by the superior performance of RUM on chemical datasets, we plan to apply our method to the equivariant modeling of $n$-body physical systems [70].
@@ B.2 RUM attenuates over-squashing
In Table 6, we conduct a brief ablation study where we delete, one by one, the components introduced in Section 3. $\omega_{u}=\mathbf{0}$ and $\omega_{x}=\mathbf{0}$ refer to the deletion of the topological and semantic representations of walks, respectively. Neglecting topological information results in a moderate decrease in performance, whereas neglecting semantic representation is more detrimental. The $\omega_{u}=\mathbf{0}$ also resembles Jin et al. [22] albeit with different walk-wise aggregation. This offers a qualitative comparison between our work and Jin et al. [22] as no source code was released for this package so no rigorous comparison was possible. We also see that the regularization methods are helpful to the performance, with self-supervision being more crucial. We attribute this effect to firstly the relative simplicity of the Cora classification task, and secondly the flexibility of the (overparametrized) RNNs.

### v3

@@ 1 Introduction: Convolutions in GNNs
Graph neural networks (GNNs) [1, 2, 3, 4, 5]—neural models operating on representations of nodes ($\mathcal{V}$) and edges ($\mathcal{E}$) in a graph (denoted by $\mathcal{G}=\{\mathcal{V},\mathcal{E}\},\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$, with structure represented by the adjacency matrix $A_{ij}=\mathbbm{1}[(v_{i},v_{j})\in\mathcal{E}]$)—have shown promises in a wide array of social and physical modeling applications. Most GNNs follow a convolutional scheme, where the $D$-dimensional node representations $\mathbf{X}\in\mathbb{R}^{|V|\times D}$ are aggregated based on the structure of local neighborhoods:
@@
In this paper, we propose a variant of GNN that does not engage the convolution operator (Equation 1) at all. [^1] Code at: https://github.com/yuanqing-wang/rum/tree/main Specifically, we stochastically sample a random walk terminating at each node and use the anonymous experiment associated with the random walk to describe its topological environment. This is combined with the semantic representations along the walk and fed into a recurrent neural network layer [16] to form the node embedding, which we call the unifying memory. We theoretically (§ 4) and experimentally (§ 5) show that the resulting model, termed random walk with unifying memory (RUM) relieves the aforementioned symptoms and offers a compelling alternative to the popular convolution-based GNNs.
@@ 2 Related works: ways to walk on a graph
More importantly, focusing mostly on expressiveness, no aforementioned walk-based or random walk kernel-based GNNs address the issue of over-smoothing and over-squashing in GNNs. Some of these works are also de facto convolutional, as the random walks are only incorporated as features in the message-passing scheme. Interestingly, most of these works are either not applicable to, or have not been tested on, node-level tasks. In the experimental § 5, we show that RUM not only outperforms these baselines on most graph-level tasks (Table 3) but also competes with a wide array of state-of-the-art convolutional GNNs on node-level tests (Table 2). Moreover, random walk kernel-based architectures, which explicitly enumerate random walk-generated subgraphs are typically much slower than WL-equivalent convolutional GNNs, whereas RUM is faster than even the simplest variation of convolutional GNN (Figure 4).
@@ 3 Architecture: combining topologic and semantic trajectories of walks
where $D(v_{i})=\sum A_{ij}$ is the degree of the node $v_{i}$. Walks originating from or terminating at any given node $v$ can thus be easily generated using this Markov chain. We record the trajectory of embeddings associated with the walk as $\omega_{x}(w)=(\mathbf{X}_{i})=(\mathbf{X}_{0},\mathbf{X}_{1},\ldots,\mathbf{X}_{l}).$ In this paper, we only consider finite-long $l$-step random walk $|w|=l\in\mathbb{Z}^{+}$. In our implementation, the random walks are sampled ad hoc during each training and inference step directly on GPU using Deep Graph Library [36] (see Appendix § A). Moreover, the walk considered here is not necessarily a path, as repeated traversal of the same node $v_{i}=v_{j},i\neq j$ is not only permitted, but also crucial to effective topological representation, as discussed below.
@@
Obvious choices to model $f$ include a feed-forward neural network after concatenation, and $\phi_{x},\phi_{u}$ recurrent neural networks (RNNs). This implies that, different from most convolutional GNNs, parameter sharing is natural and the number of parameters is going to stay constant as the model incorporates a larger neighborhood. Compared to dot product-based, transformer-like modules [38], RNNs not only have linear complexity (see detailed discussions below) w.r.t. the sequence length but also naturally encodes the inductive bias that nodes closer to the origin have stronger impact on the representation. The gated recurrent unit (GRU)[16] variant is used everywhere in this paper. Additional regularizations are described in Appendix § B.1.
@@
Assumption 2 is easy to satisfy for feed-forward neural networks [40] and RNNs [41]. Composing injective functions, we remark that $h(w)$ is also injective w.r.t. $\omega_{x}(w)$ and $\omega_{u}(w)$; despite of Assumption 3, our analysis can be extended to disjointed graphs by restricting the analysis to the connected regions in a graph. Under such assumptions, we show, in Remark 8 (deferred to the Appendix), that $\psi$ is injective, meaning that nodes with different random walks will have different distributions of representations $\psi(v_{1})\neq\psi(v_{2})$. We also refer the readers to the Theorem 1 in Micali and Zhu [37] for a discussion on the representation power of anonymous experiments on unlabelled graphs. Combining with the semantic representations and promoting the argument from a node level to a graph level, we arrive at:
@@
The main idea of the proof of Theorem 4 (in Appendix § D.2) involves explicitly enumerating all possible non-isomorphic structures for graphs with 3 nodes and showing, by induction, that if the theorem stands for graph of $N-1$ size it also holds for $N$-sized graphs. We also show in Appendix § D.1 that a number of key graph properties such as cycle size (Example 8.1) and radius (Example 8.2) that convolutional GNNs struggle [43, 44] to learn can be analytically expressed using $\omega_{u}$. As these are solely functions of $\omega_{x}$, they can be approximated arbitrarily well by universal approximators. These examples are special cases of the finding that RUM is stricly more expressive than Weisfeiler-Lehman isomorphism test [9]:
@@ 4.2 RUM alleviates over-smoothing and over-squashing
is deferred to Appenxix § B.2, where we verify that the inter-node Jacobian $|\partial\mathbf{X}_{v}^{(l+1)}/\partial\mathbf{X}_{u}^{(0)}|$ decays slower as the distance between $u,v$ grows vis-à-vis the convolutional counterparts. Briefly, although RUM does not address the information bottleneck with exponentially growing reception field (the $1/(\hat{A}^{l+1})_{uv}$ term in Equation 16), it nevertheless can have a non-vanishing (nor exploding) gradient from the aggregation function ($|\nabla\phi_{x}|$).
@@ 5 Experiments
On a wide array of real-world node- and graph-level tasks, we benchmark the performance of RUM to show its utility in social and physical modeling. Next, to thoroughly examine the performance of RUM, we challenge it with carefully designed illustrative experiments. Specifically, we ask the following questions in this section, with Q1, Q2, and Q3 already theoretically answered in § 4: Q1: Is RUM more expressive than convolutional GNNs? Q2: Does RUM alleviate over-smoothing? Q3: Does RUM alleviate over-squashing? Q4: Is RUM slower with convolutional GNNs? Q5: Is RUM robust? Q6: How does RUM scale up to huge graphs? Q7: What components of RUM are contributing most to the performance of RUM?
@@
Having illustrated in § 4 that RUM can distinguish non-isomorphic graphs, we experimentally test this insight on the popular Circular Skip Link dataset [58, 59]. Containing 4-regular graph with edges connected to form a cycle and containing skip-links between nodes, this dataset is artificially synthesized in Murphy et al. [58] to create an especially challenging task for GNNs. As shown in Appendix Table 7, all convolutional GNNs fail to perform better than a constant baseline (there are 10 classes uniformly distributed). 3WLGNN [60], a higher-order GNN of at least $\mathcal{O}(2)$ complexity that operates on explicitly enumerated triplets of graphs, can distinguish these highly similar 4-regular graphs by comparing subgraphs. RUM, with $\mathcal{O}(|\mathcal{N}|)$ linear complexity, achieves similarly high accuracy. One can think of RUM as a stochastic approximation of the higher-order GNN, with all of its explicitly enumerated subgraphs being identified by RUM with a probability that decreases with the complexity of the subgraph.
@@
Having studied the relationship between inference speed and the walk lengths and number of samples, we furthermore study its impact on performance. Using Cora [45] citation graph and vary the walk lengths and number of samples from $1$ to $9$, where the performance of RUM improves as more samples are taken and longer walks are employed, though more than 4 samples and walks longer than $L>4$ yield qualitatively satisfactory results; this empirical finding has guided our hyperparameter design. In Figure 2, we also compare the Dirichlet energy of RUM-generated layer representations with not only baselines GCN [1] and GAT [6], but also strategies to alleviate over-smoothing discussed in § 2, namely residual connection and stochastic regularization [35, 26, 61], and show that when $L$ gets large, only RUM can maintain Dirichlet energy. Traditionally, since Kipf and Welling [1] (see its Figure 5 compared to Figure 4), the best performance on Cora graph was found with $2$ or $3$ message-passing rounds, since mostly local interactions are dominating the classification, and more rounds of message-passing almost always lead to worse performance. As theoretically demonstrated in § 4.2, RUM is not as affected by these symptoms. Thus, RUM is especially appropriate for modeling long-range interactions in graphs without sacrificing local representation power.
@@
To verify that RUM indeed alleviates over-squashing (§ B.2), in Figure 6, we adopt the tree neighborhood matching synthetic task in Alon and Yahav [14] where binary tree graphs are proposed with the label and the attributes of the root matching a faraway leave. The full discussion is moved to the Appendix § B.3.
@@
Though both have linear runtime complexity (See § 3), intuitively, it might seem that RUM would be slower than convolution-based GCN due to the requirement of multiple random walk samples. This is indeed true for CPU implementations shown in Figure 4 left. When equipped with GPUs (specs in Appendix § A), however, RUM is significantly faster than even the simplest convolutional GNN—GCN [1]. It is worth mentioning that the GCN and GAT [6] results were harvested using the heavily optimized Deep Graph Library [36] sparse implementation whereas RUM is implemented naïvely in PyTorch [62], though the popular GRU component [16] have already undergone CUDA-specific optimization.
@@ 6 Conclusions
We design an innovative GNN that uses an RNN to unify the semantic and topological representations along stochastically sampled random walks, termed random walk with unifying memory (RUM) neural networks. Free of the convolutional operators, our methodology does not suffer from symptoms characteristic of Laplacian smoothing, including limited expressiveness, over-smoothing, and over-squashing. Most notably, our method is more expressive than the Weisfeiler-Lehman isomorphism test and can distinguish all non-isomorphic graphs up to the reconstruction conjecture. Thus, it is more expressive than all of convolutional GNNs equivalent to the WL test, as we demonstrate theoretically in § 4. RUM is significantly faster on GPUs than even the simplest convolutional GNNs (§ 5) and shows superior performance across a plethora of node- and graph-level benchmarks.
@@
Very dense graphs. As evidenced by the underwhelming performance of the Computer [49] dataset and discussed in § 5, RUM might suffer from high variance with very dense graphs (average degree over 15). Tottering. In our implementation, we have not ruled out the 2-cycles from the random walks, as that would require specialized implementation for walk generation. This, however, would reduce the average information content in a fixed-length random walk (known as tottering [64]). We plan to characterize the effect of excluding these walks. Biased walk. Here, we have only considered unbiased random walk, whereas biased random walk might display more intriguing properties as they tend to explore faraway neighborhoods more effectively [17]. Directed graphs. Since we have only developed RUM for undirected graph (random walk up to a random length is not guaranteed to exist for directed graphs), we would have to symmetrize the graph to perform on directed graphs (such as the heterophilic datasets [50]); this create additional information loss and complexity.
@@
Theoretical. We plan to expand our theoretical framework to account for the change in layer width and depth to derive analytical estimates for realizing key graph properties. Applications. Random walks are intrinsically applicable to uncertainty-aware learning. We plan to incorporate the uncertainty naturally afforded by the model to design active learning models. On the other hand, the physics-based graph modeling field is also heavily dominated by convolutional GNNs. Inspired by the superior performance of RUM on chemical datasets, we plan to apply our method in drug discovery settings [65, 66, 67, 68, 69] and furthermore on the equivariant modeling of $n$-body physical systems [70].
@@ B.2 RUM attenuates over-squashing
In Table 6, we conduct a brief ablation study where we delete, one by one, the components introduced in § 3. $\omega_{u}=\mathbf{0}$ and $\omega_{x}=\mathbf{0}$ refer to the deletion of the topological and semantic representations of walks, respectively. Neglecting topological information results in a moderate decrease in performance, whereas neglecting semantic representation is more detrimental. The $\omega_{u}=\mathbf{0}$ also resembles Jin et al. [17] albeit with different walk-wise aggregation. This offers a qualitative comparison between our work and Jin et al. [17] as no source code was released for this package so no rigorous comparison was possible. We also see that the regularization methods are helpful to the performance, with self-supervision being more crucial. We attribute this effect to firstly the relative simplicity of the Cora classification task, and secondly the flexibility of the (overparametrized) RNNs.